# Personalized Topology-Informed 12-Lead ECG Electrode Localization from Incomplete Cardiac MRIs for Efficient Cardiac Digital Twins

### v1

@@ Abstract
Cardiac digital twins (CDTs) offer personalized in-silico cardiac representations for the inference of multi-scale properties tied to cardiac mechanisms. The creation of CDTs requires precise information about the electrode position on the torso, especially for the personalized electrocardiogram (ECG) calibration. However, current studies commonly rely on additional acquisition of torso imaging and manual/semi-automatic methods for ECG electrode localization. In this study, we propose a novel and efficient topology-informed model to fully automatically extract personalized ECG electrode locations from 2D clinically standard cardiac MRIs. Specifically, we obtain the sparse torso contours from the cardiac MRIs and then localize the electrodes from the contours. Cardiac MRIs aim at imaging of the heart instead of the torso, leading to incomplete torso geometry within the imaging. To tackle the missing topology, we incorporate the electrodes as a subset of the keypoints, which can be explicitly aligned with the 3D torso topology. The experimental results demonstrate that the proposed model outperforms the time-consuming conventional method in terms of accuracy (Euclidean distance: $1.24\pm 0.293$ cm vs. $1.48\pm 0.362$ cm) and efficiency ($2$ s vs. $30$-$35$ min). We further demonstrate the effectiveness of using the detected electrodes for in-silico ECG simulation, highlighting their potential for creating accurate and efficient CDT models. The code will be released publicly after the manuscript is accepted for publication.
@@
Cardiac digital twins , 12-lead ECG , Cardiac MRI , Electrodes , Topology-informed model , Incomplete data
@@ 1 Introduction
Figure 1: Visualization of the setup of the torso with 10 electrodes (labeled with cyan-blue dots) and 12-lead clinical ECG recorded on the torso.
@@
Cardiac digital twins (CDTs), also known as virtual heart models, aim to replicate the anatomical and physiological characteristics of an individual heart (Corral-Acero et al., 2020; Gillette et al., 2021; Camps et al., 2024b). Their ability to simulate and predict personalized responses to interventions makes them invaluable tools for advancing precision medicine and improving patient outcomes in cardiovascular care (Arevalo et al., 2016; Boyle et al., 2019; Prakosa et al., 2018). The development of CDTs becomes possible due to the availability of patient-specific information from electrical data (e.g. electrocardiogram, ECG) and imaging data (Li et al., 2024b). ECG, a widely utilized tool for recording the electrical activity within the heart, is commonly performed with 12 leads. The 12-lead ECG utilizes 10 electrodes (see Fig. 1) to measure body surface potential, inferring cardiac potential based on heart and torso anatomy and their relative positions (Roy et al., 2020; Gillette et al., 2022). The variations in electrode positions can influence the duration and amplitudes of recorded ECG waveforms (Kania et al., 2014; Rjoob et al., 2020). Therefore, the electrode positions serve as a key factor in accurately capturing the relationship between the electrical activity of the heart and the surrounding torso anatomy (van Dam et al., 2014). However, due to the unavailability of precise electrode positions and torso imaging in clinical practice, many studies use standardized ECG electrode locations for CDT, neglecting actual anatomical variations in patients (van Dam et al., 2014; Qian et al., 2023).
@@
Cardiac imaging, typically obtained through magnetic resonance imaging (MRI) or computed tomography (CT) scans, offers detailed insights into the anatomical, morphological, and functional features of the heart. These imaging techniques have proved invaluable for reconstructing cardiac anatomy and pinpointing pathology locations (Li et al., 2023b). However, standard cardiac imaging poses challenges in reconstructing torso geometry and localizing electrodes (Smith et al., 2023, 2022). Since these imaging views primarily focus on the heart, there is limited information available on the contours of the sections of the torso distant from the heart, leading to an incomplete representation of the torso, as presented in Fig. 2. Furthermore, unwanted features such as the head, neck, and arms may be present in the images, which can hinder and negatively impact torso reconstruction (Gillette et al., 2015). Also, common MRI artifacts often manifest as shadow regions, particularly above the shoulder or near the waist.
@@
Therefore, additional torso images are usually required for heart-torso geometry modeling and electrode localization (Ramanathan et al., 2004; Gillette et al., 2021, 2022). However, obtaining full torso geometry necessitates additional scanning time, leading to heightened patient discomfort due to prolonged breath-holding and an escalation in the cost of MRI data acquisition. Few studies attempted to reconstruct torso geometry directly from cardiac MRI and then localize the electrodes, which was challenging due to partial torso shape information (Zacur et al., 2017; Smith et al., 2022). These studies commonly integrated statistical shape models (SSMs) into the reconstruction process to impose prior anatomical constraints about the torso geometry, which however were often time-consuming (Smith et al., 2023, 2022).
@@
In this work, we develop a novel pipeline to localize the ECG electrodes from 2D clinically standard cardiac MRIs. The proposed method incorporates topology information about the torso into the pipeline to guide the electrode localization directly from incomplete contours, which are extracted from cardiac MRIs. This is achieved by introducing the keypoint detection and surface skeleton-assisted point cloud completion (PCN) modules. To learn the spatial relationship between electrodes and torso geometry, we adopt the electrodes as a subset of the keypoints, which are then used for torso reconstruction. Therefore, an end-to-end learning framework is created for ECG electrode localization and 3D torso reconstruction from incomplete contours. To the best of our knowledge, this is the first fully automatic deep learning based 12-lead ECG electrode localization work, addressing the challenges of incomplete information through the incorporation of topology information. The main contributions of this work include:
@@
    We develop a novel topology-informed method for estimating 12-lead ECG electrode location directly from incomplete torso contours.
@@
Figure 2: Visualization of extracted full and refined contours and the 3D arrangement of the refined torso contours (only partially visualized here).
@@ 2.1 Cardiac Digital Twins: Bridging Precision Cardiology
CDT technology aims to create a personalized virtual cardiac representations based on patient-specific data, such as imaging, ECG, etc (Li et al., 2024b; Camps et al., 2024a). It has shown great promise in the personalized treatments for cardiac disease, such as the stratification of arrhythmia risk (Arevalo et al., 2016) and the ablation guidance of persistent atrial fibrillation (Boyle et al., 2019). The creation of a digital heart usually involves two stages, namely anatomical twinning and functional twinning (Gillette et al., 2021; Li et al., 2024b). For the anatomical twining, one needs to perform the cardiac segmentation, 3D heart-torso geometry reconstruction, and the modeling of relevant pathology. This process is complicated by the variability in cardiac anatomy among different individuals and by imaging artifacts and noise. The functional twining mainly includes the ECG simulation and calibration via solving the ECG inverse problem, which is inherently ill-posed (Li et al., 2024a). This process is further complicated by the limitations of ECG recordings, which are sparse, noisy, and subject to substantial uncertainties. Furthermore, the torso geometry and electrode localization accuracy can influence the inverse solution, as the electrophysiology measurement is generally performed on the body surface (Zemzemi et al., 2015; Qian et al., 2023).
2.2 Torso Geometry Reconstruction and Electrode Localization
Most existing electrode localization works rely on the accurate 3D torso reconstruction from medical imaging (such as CT, MRI, fluoroscopy, etc) (Mincholé et al., 2019; Gillette et al., 2021; Salvador et al., 2024). They generally are based on manual or semi-automatic methods and require additional torso imaging techniques (Mincholé et al., 2019; Gillette et al., 2021). For example, Ramanathan et al. (2004) manually segmented slice by slice to obtain heart geometry (as epicardial contours on each slice) and torso geometry (described by body-surface electrode positions) from CT images. The CT images are often combined with X-ray-opaque Ag/AgCl electrodes, allowing for simultaneous acquisition of body-surface electrode locations. Gillette et al. (2021) segmented the torso from the full torso MRI using threshold-based filters with boolean operations in Seg3D [^1] https://www.sci.utah.edu/cibc-software/seg3d.html and then generated anatomical torso meshes using an image-based unstructured mesh generation technique. Salvador et al. (2024) created torso models from CT images using semi-automatic methods, i.e., threshold- and region-growing-based segmentation methods.
@@ 2.2 Torso Geometry Reconstruction and Electrode Localization
Instead of acquiring torso imaging, several studies reconstructed torso geometry from cardiac MRI and then localized the electrodes on it (Zacur et al., 2017; Smith et al., 2022). To consider the torso heterogeneity, Zacur et al. (2017) expanded the torso representation to encompass ribs and lungs and then virtually placed electrodes on it to synthesize a 12-lead ECG. Nevertheless, including various organs in computational studies remains controversial, as some research indicate minimal contributions of organ impedances on body surface potential (BSP) (Geneser et al., 2008), while others emphasize significant effects attributable to uncertainties in impedance values (Keller et al., 2010). Instead of relying on torso reconstruction, few studies directly employed photography for electrode localization (Ghanem et al., 2003; Perez-Alday et al., 2018; Roudijk et al., 2021). For example, Ghanem et al. (Ghanem et al., 2003) employed 2D photography (SONY DSC-S85) and point reconstruction techniques to determine ECG electrode positions, calibrating with a known object and matching 3D coordinates of identified electrodes in sequential photographs. Nowadays, 3D cameras such as Intel Real Sense D435 can be employed to directly capture 3D photos, which can be fitted to a standard thoracic model and then the electrode positions can be segmented based on these 3D photos (Roudijk et al., 2021).
@@
Figure 3: Topology-informed model (TIM) for 12-lead ECG electrode detection and torso geometry reconstruction from sparse and incomplete torso contours. Here, the electrodes are labeled via red dots within the predicted keypoints, which are then converted into surface skeleton to guide the 3D torso reconstruction. Note that the diagram takes 32 keypoints as an example. PCN: point completion model.
@@ 2.3 Incomplete Data Processing
In real-world applications, the data may be partially missing, corrupted by noise, or exhibit other forms of incompleteness (Aste et al., 2015). In the context of medical imaging, incomplete data refers to scenarios where the images lack coverage of specific regions of interest (ROIs), or when the dataset of the target patient includes a series of imaging sequences or views, with some of these sequences or views being absent. It may arise due to various factors such as imaging limitations, out-of-range measurements, or challenges in capturing specific ROIs. For example, many head CT images in clinical practice are obtained without conducting a complete head scan, aiming to minimize radiation exposure (Wan et al., 2023). The subjects with myocardial infarction may not have full sequences for the pathology analysis, such as missing late gadolinium enhancement (LGE) MRI or mapping MRI scans, due to patient-specific conditions (Qiu et al., 2023). The long-axis MRI data may be missing for 3D cardiac geometry reconstruction, resulting in incomplete shape information especially at the apex and base of the heart (Sander et al., 2023). Therefore, it is essential to develop models with the capability of managing incomplete imaging information in a flexible manner. Traditional methods of processing incomplete imaging data often involve incorporating prior knowledge, such as SSM, to help reconstruct or complete missing information (Zacur et al., 2017; Smith et al., 2022). These methods require expensive optimizations during inference and thus are impractical for real-time applications. To mitigate this issue, Wan et al. (2023) embedded the SSM into the model training to guide the learning process for solutions that align with the prior knowledge. Furthermore, data-driven image synthesis has been widely used for processing incomplete imaging data (Wang et al., 2020; Zhang et al., 2021). Integrating information from non-imaging data can be a valuable strategy for completing missing knowledge within imaging data (Li et al., 2024b). Recently, several works used a deep learning based parameterized model to directly map partial input to a complete shape (Yuan et al., 2018; Tang et al., 2022; Beetz et al., 2023). However, these approaches do not explicitly learn the topology of the complete shape, resulting in less robust predictions, particularly in missing regions.
@@ 3.1 Torso Contour Extraction from Multi-View Cardiac MRIs
We use a two-step automated framework that comprises a segmentation network followed by a two-channel refinement network for the torso contour extraction (Smith et al., 2022). Specifically, we firstly employ a U-Net based torso segmentation model to extract all visible sections of the body and then convert it into full contours. Next, a U-Net based refinement model is utilized to eliminate contour sections influenced by undesired anatomical features (e.g., head, neck, and connected arm sections) and any boundaries caused by MRI artifacts. Note that the input of the refinement model include both the image and the corresponding full contour obtained in the first step. Both models are optimized by maximizing the Dice score between the predicted mask and ground truth mask. As shown in Fig. 2, the extracted torso contours can be arranged together into a sparse torso representation based on their 3D world coordinates. Nonetheless, the torso representation is sparse and incomplete, and thus some of the electrodes may not be described in the sparse contours.
Figure 4: Statistical shape model (SSM) based 3D torso mesh reconstruction and electrode localization from multi-view 2D cardiac MRIs.

### v2

# Personalized Topology-Informed Localization of Standard 12-Lead ECG Electrode Placement from Incomplete Cardiac MRIs for Efficient Cardiac Digital Twins

@@ Abstract
Cardiac digital twins (CDTs) offer personalized in-silico cardiac representations for the inference of multi-scale properties tied to cardiac mechanisms. The creation of CDTs requires precise information about the electrode position on the torso, especially for the personalized electrocardiogram (ECG) calibration. However, current studies commonly rely on additional acquisition of torso imaging and manual/semi-automatic methods for ECG electrode localization. In this study, we propose a novel and efficient topology-informed model to fully automatically extract personalized ECG standard electrode locations from 2D clinically standard cardiac MRIs. Specifically, we obtain the sparse torso contours from the cardiac MRIs and then localize the standard electrodes of 12-lead ECG from the contours. Cardiac MRIs aim at imaging of the heart instead of the torso, leading to incomplete torso geometry within the imaging. To tackle the missing topology, we incorporate the electrodes as a subset of the keypoints, which can be explicitly aligned with the 3D torso topology. The experimental results demonstrate that the proposed model outperforms the time-consuming conventional model projection-based method in terms of accuracy (Euclidean distance: $1.24\pm 0.293$ cm vs. $1.48\pm 0.362$ cm) and efficiency ($2$ s vs. $30$-$35$ min). We further demonstrate the effectiveness of using the detected electrodes for in-silico ECG simulation, highlighting their potential for creating accurate and efficient CDT models. The code is available at https://github.com/lileitech/12lead_ECG_electrode_localizer.
@@
[…]
@@ 1 Introduction
Fig. 1: Illustration of the setup of the torso with electrodes (labeled with cyan-blue dots) and simulated 12-lead ECG.
@@
Cardiac digital twins (CDTs), also known as virtual heart models, aim to replicate the anatomical and physiological characteristics of an individual heart [14, 9]. Their ability to simulate and predict personalized responses to interventions makes them invaluable tools for advancing precision medicine and improving patient outcomes in cardiovascular care [1, 6, 31]. The development of CDTs becomes possible due to the availability of patient-specific information from electrical data (e.g. electrocardiogram, ECG) and imaging data [22]. ECG, a widely utilized tool for recording the electrical activity within the heart, is commonly performed with 12 leads. The standard 12-lead ECG utilizes 10 electrodes to measure body surface potential, inferring cardiac potential based on heart and torso anatomy and their relative positions [38, 15]. The variations in electrode positions can influence the duration and amplitudes of recorded ECG waveforms [17, 36]. Therefore, the electrode positions serve as a key factor in accurately capturing the relationship between the electrical activity of the heart and the surrounding torso anatomy [10]. Accurate electrode positioning typically requires specialized devices, such as custom ECG systems with electrode localizers or specialized photography, which are generally unavailable in routine clinical practice [14, 12, 37]. Furthermore, given the challenges of localizing standard limb electrodes via medical imaging, the Mason-Likar (M-L) modification of the standard 12-lead ECG has been developed. It relocates limb electrodes (RA, LA, RL, LL) from the arms and legs to the torso while retaining the precordial leads (V1–V6) in their standard chest positions [25], as shown in Fig. 1. Although this altered placement may alter ECG morphology, a study by Man et al. [25] demonstrated that standard 12-lead ECGs can be effectively reconstructed from M-L configurations using specific transformation matrices. Consequently, current studies on creating digital twins primarily employ standard torso-based M-L electrode locations identified through medical imaging [8, 24, 15, 32, 9, 7, 39, 22].
@@
Cardiac imaging, typically obtained through magnetic resonance imaging (MRI) or computed tomography (CT) scans, offers detailed insights into the anatomical, morphological, and functional features of the heart [23]. These imaging techniques have proved invaluable for reconstructing cardiac anatomy and pinpointing pathology locations. However, standard cardiac imaging poses challenges in reconstructing torso geometry and localizing electrodes [42, 41]. Since these imaging views primarily focus on the heart, there is limited information available on the contours of the sections of the torso distant from the heart, leading to an incomplete representation of the torso, as presented in Fig. 2. Furthermore, unwanted features such as the head, neck, and arms may be present in the images, which can hinder and negatively impact torso reconstruction [16]. Also, common MRI artifacts often manifest as shadow regions, particularly above the shoulder or near the waist.
@@
Therefore, additional torso images are usually required for heart-torso geometry modeling and electrode localization [35, 14, 15]. However, obtaining full torso geometry necessitates additional scanning time, leading to heightened patient discomfort due to prolonged breath-holding and an escalation in the cost of MRI data acquisition. Few studies attempted to reconstruct torso geometry directly from cardiac MRI and then localize the electrodes, which was challenging due to partial torso shape information [48, 41]. These studies commonly integrated statistical shape models (SSMs) into the reconstruction process to impose prior anatomical constraints about the torso geometry, which however were often time-consuming. This is because SSM-based electrode detection typically involves iterative optimization of 3D torso geometry by aligning the individual shape with a mean torso model.
@@
In this work, we develop a novel pipeline to efficiently localize the standard torso electrodes from 2D clinically acquired cardiac MRIs. Note that the cardiac imaging setup includes scout images that partially cover the torso to serve as localizers for the subsequent cardiac cine imaging. The proposed method incorporates topology information about the torso into the pipeline to guide the electrode localization directly from incomplete contours, which are extracted from cardiac MRIs. This is achieved by introducing the keypoint detection and surface skeleton-assisted point cloud completion (PCN) modules. As shown in Fig. 1, ECG electrodes are positioned at anatomically significant locations to capture cardiac electrical activity, particularly the limb electrodes, which serve as key landmarks for the torso geometry (specifically the front side of the torso). To employ the spatial relationship between the electrodes and torso geometry, we adopt the electrodes as a subset of the keypoints used for torso reconstruction. Therefore, an end-to-end learning framework is created for ECG electrode localization and 3D torso reconstruction from incomplete contours. To the best of our knowledge, this is the first fully automatic deep learning based torso electrode localization work, addressing the challenges of incomplete information through the incorporation of topology information. The main contributions of this work include:
@@
    We develop a novel topology-informed method for estimating standard torso electrode location from 2D clinically acquired cardiac MRIs.
@@
Fig. 2: Visualization of extracted full and refined contours from cardiac MRIs and the 3D arrangement of the refined torso contours (only partially visualized here). Here, the acquired cardiac MRIs include scout images covering a larger range of the torso, typically used as localizers, and heart-focused slices in various views, including long-axis (LAX) views (2-, 3-, and 4-chamber) and a stack of short-axis (SAX) views.
@@ 2.1 Cardiac Digital Twins: Bridging Precision Cardiology
CDT technology aims to create a personalized virtual cardiac representations based on patient-specific data, such as imaging, ECG, etc [22, 7]. It has shown great promise in the personalized treatments for cardiac disease, such as the stratification of arrhythmia risk [1] and the ablation guidance of persistent atrial fibrillation [6]. The creation of a digital heart usually involves two stages, namely anatomical twinning and functional twinning [14, 22]. For the anatomical twining, one needs to perform the cardiac segmentation, 3D heart-torso geometry reconstruction, and the modeling of relevant pathology. This process is complicated by the variability in cardiac anatomy among different individuals and by imaging artifacts and noise. The functional twining mainly includes the ECG simulation and calibration via solving the ECG inverse problem, which is inherently ill-posed [20]. This process is further complicated by the limitations of ECG recordings, which are sparse, noisy, and subject to substantial uncertainties. Furthermore, the torso geometry and electrode localization accuracy can influence the inverse solution, as the electrophysiological (EP) measurement is generally performed on the body surface [50, 32].
2.2 Torso Reconstruction and Electrode Localization
@@
Most existing electrode localization studies focus on identifying standard M-L electrode positions on the torso for ECG simulation [26, 14, 39]. These methods typically rely on semi-automatic techniques and often require additional torso imaging for precise 3D torso reconstruction [26, 14]. For example, Salvador et al. [39] created heart-torso models from full torso CT scan using semi-automatic methods, i.e., threshold- and region-growing-based segmentation methods, and then electrodes were identified on the torso. Giffard-Roisin et al. [13] reconstructed the torso mesh from CT images and identified the electrodes for 12-lead ECG from body surface potential (BSP) mapping electrodes based on the standard torso ECG placement. Instead of acquiring torso imaging, several studies reconstructed torso geometry from cardiac MRIs and then localized the standard torso electrodes on the 3D torso [48, 41]. To consider the torso heterogeneity, Zacur et al. [48] expanded the torso representation to encompass ribs and lungs and then virtually placed electrodes on it to synthesize a 12-lead ECG. Nevertheless, including various organs in computational studies remains controversial, as some research indicates minimal contributions of organ impedances on BSP [11], while others emphasize significant effects attributable to uncertainties in impedance values [18].
Instead of localizing the virtual electrodes on the torso, a few studies directly employed specialized CT/ MRI or photography to localize the actual electrodes [12, 28, 37]. For example, Ramanathan et al. [35] manually segmented slice by slice to obtain heart-torso geometry from CT images, which are combined with X-ray-opaque Ag/AgCl electrodes, allowing for simultaneous acquisition of body-surface electrode locations. Gillette et al. [14] employed MRI-compatible electrodes that were placed before image acquisition to obtain the electrode location. Then, they segmented the torso from the full torso MRI using threshold-based filters with boolean operations in Seg3D [^1] https://www.sci.utah.edu/cibc-software/seg3d.html and then generated anatomical torso meshes using an image-based unstructured mesh generation technique. Ghanem et al. [12] employed 2D photography (SONY DSC-S85) and point reconstruction techniques to determine ECG electrode positions, calibrating with a known object and matching 3D coordinates of identified electrodes in sequential photographs. Nowadays, 3D cameras such as Intel Real Sense D435 can be employed to directly capture 3D photos, which can be fitted to a standard thoracic model and then the electrode positions can be segmented based on these 3D photos [37]. While these approaches utilize specialized imaging for precise electrode localization of the actual ECG electrode placement, our work focuses on detecting the standard electrode positions of 12-lead ECG using routine available imaging, addressing a distinct challenge in clinical practice.
@@
Fig. 3: Topology-informed model (TIM) for 12-lead ECG electrode detection and torso geometry reconstruction from sparse and incomplete torso contours. Here, the electrodes are labeled via red dots within the predicted keypoints, which are then converted into surface skeleton to guide the 3D torso reconstruction. Note that the diagram takes 32 keypoints as an example. PCN: point completion model.
@@ 2.3 Incomplete Data Processing
In real-world applications, the data may be partially missing, corrupted by noise, or exhibit other forms of incompleteness [2]. In the context of medical imaging, incomplete data refers to scenarios where the images lack coverage of specific regions of interest (ROIs), or when the dataset of the target patient includes a series of imaging sequences or views, with some of these sequences or views being absent. It may arise due to various factors such as imaging limitations, out-of-range measurements, or challenges in capturing specific ROIs. For example, many head CT images in clinical practice are obtained without conducting a complete head scan, aiming to minimize radiation exposure [45]. The subjects with myocardial infarction may not have full sequences for the pathology analysis, such as missing late gadolinium enhancement (LGE) MRI or mapping MRI scans, due to patient-specific conditions [33]. The long-axis MRI data may be missing for 3D cardiac geometry reconstruction, resulting in incomplete shape information especially at the apex and base of the heart [40]. Therefore, it is essential to develop models with the capability of managing incomplete imaging information in a flexible manner. Traditional methods of processing incomplete imaging data often involve incorporating prior knowledge, such as SSM, to help reconstruct or complete missing information [48, 41]. These methods require expensive optimizations during inference and thus are impractical for real-time applications. To mitigate this issue, Wan et al. [45] embedded the SSM into the model training to guide the learning process for solutions that align with the prior knowledge. Furthermore, data-driven image synthesis has been widely used for processing incomplete imaging data [46, 51]. Integrating information from non-imaging data can be a valuable strategy for completing missing knowledge within imaging data [22]. Recently, several works used a deep learning based parameterized model to directly map partial input to a complete shape [47, 43, 4]. However, these approaches do not explicitly learn the topology of the complete shape, resulting in less robust predictions, particularly in missing regions.
@@ 3.1 Torso Contour Extraction from Multi-View Cardiac MRIs
The cardiac MRIs are acquired using standard clinical protocols not designed for torso detection, often including structures and artifacts that interfere with reconstruction, and sometimes missing parts of the region of interest. The reader is referred to Sec. 4.1 for detailed information on the MRI data used in the experiments. Therefore, we use a two-step automated framework that comprises a segmentation network followed by a two-channel refinement network for the torso contour extraction [41]. Specifically, we firstly employ a U-Net based torso segmentation model to extract all visible sections of the body and then convert it into full contours. Next, a U-Net based refinement model is utilized to eliminate contour sections influenced by undesired anatomical features (e.g., head, neck, and connected arm sections) and any boundaries caused by MRI artifacts. Note that the input of the refinement model include both the image and the corresponding full contour obtained in the first step. Both models are optimized by maximizing the Dice score between the predicted mask and ground truth mask. As shown in Fig. 2, the extracted torso contours can be arranged together into a sparse torso representation based on their 3D world coordinates. Nonetheless, the torso representation is sparse and incomplete, and thus some of the electrodes may not be described in the sparse contours.
Fig. 4: Statistical shape model (SSM) based 3D torso mesh reconstruction and electrode localization from multi-view 2D cardiac MRIs.